OpenStudy (anonymous):

A first order linear equation in the form y'+p(x)y=f(x) can be solved by finding an integrating factor mu(x) = exp( int (p(x) dx)) (1) Given the equation y' + 3 y = 2 find mu(x) = (2) Then find an explicit general solution with arbitrary constant C. y = (3) Then solve the initial value problem with y(0)=2 y =

OpenStudy (anonymous):

\[y'+3y=2~~\implies~~\mu(x)=e^{\int 3~dx}=e^{3x}\] \[\begin{align*}e^{3x}y'+3e^{3x}y=2e^{3x}~~\iff~~\frac{d}{dx}\left[e^{3x}y\right]&=2e^{3x}\\ e^{3x}y&=2\int e^{3x}~dx\\ e^{3x}y&=\frac{2}{3}e^{3x}+C\\ y&=\frac{2}{3}+Ce^{-3x} \end{align*}\] \[2=\frac{2}{3}+Ce^{-3(0)}~~\iff~~\frac{4}{3}=C~~\implies~~y=\frac{2}{3}+\frac{4}{3}e^{-3x}\]